# AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels

### v1

@@ B.1. Baselines
OpenAI-Ada-002  [^21] https://openai.com/index/new-and-improved-embedding-model/.   It is a highly efficient text embedding model that converts natural language into dense vectors for a wide range of applications, including semantic search and similarity tasks.
Qwen2222222https://huggingface.co/Qwen/Qwen2-7B-Instruct.   Qwen2 is a comprehensive suite of foundational and instruction-tuned language models, encompassing a parameter range from 0.5 to 72 billion, featuring dense models and a Mixture-of-Experts model.
ChatGLM3232323https://huggingface.co/THUDM/chatglm3-6b.   ChatGLM3-6B is a next-generation conversational pre-trained model with strong performance across tasks like semantics, reasoning, and code execution, and supports complex scenarios such as tool use and function calls.
Llama2242424https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.    Llama2 is an auto-regressive language model that uses an optimized transformer architecture. The tuned versions utilize supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
B.2. Evaluation Prompt
We use the C-MTEB [^25] https://github.com/FlagOpen/FlagEmbedding/tree/master/C_MTEB framework to evaluate the performance of various retrieval models on CMIRB. To ensure stability, we set the temperature coefficient to 0 when generating hypothetical documents for LLMs, producing deterministic results. For each dataset, the prompts used to generate pseudo-documents are shown in Figure 8. The IIYIPost and CSLCite datasets utilize the T2P template to prompt LLMs to generate documents based on the given title. For the CSLRel dataset, we employ the P2P template to instruct the model to produce similar text. As for the other datasets, the Q2P template is employed by the LLM to generate answers to medical questions.
Appendix C More Experiment Results
Table 11. Prompts for Data Processing (1).
Medical Relevance Prompt (LLM.med_score)
You will receive a question-answer pair from Baidu Search. Your task is to evaluate whether the Q&A is related to the medical field and
output the result in JSON format.
The JSON object must include the following keys:
- ”reason”: a string explaining the reason for your judgment.
- ”label”: an int, 0/1.
Please adhere to the following steps:
- If the content mentioned in the question and answer includes medical information and is related to the medical field, the label should be 1.
- If most of the content in the question and answer is unrelated to the medical field, the label should be 0.
You need to make a judgment and provide a reason.
Please output the result as required, and do not output any other content.
Here is the text:
Question: [QUESTION]
Answer: [ANSWER]
Passage Reranking Prompt (LLM.reranking)
You will be given a medical question-answer pair (Q-A) and its Top-20 most relevant documents. Your task is to select up to 3 most relevant
and independent reference documents to verify or support the given answer accurately and efficiently.
Your output is a JSON object, which must contain the following keys:
- "$doc\_id"$: a list, the document IDs you selected, ranked by importance, such as "$id_{1}"$.
- ”explanation”: a string, the reasons for your selection of these documents.
Please adhere to the following steps:
- 1. Read the medical question and its answer.
- 2. Examine the Top-20 retrieved documents to identify which documents independently contain the essential information needed to
support or verify the provided answer.
- 3. If suitable documents are found, select up to 3 documents that can independently support the answer and explain why these documents
are relevant. If none of the documents are suitable, do not select any.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the medical Q-A pair and Top-20 documents:
Q-A pair: [Q-A PAIR]
Documents: [DOCUMENTS]
Evidence Extracting Prompt (LLM.extract_evidence)
You will be given a medical question, its answer and a related document. Your task is to extract evidence spans from the document that
directly or indirectly support the answer to the medical question.
Your output is a JSON object, which must contain the following keys:
- ”evidence_spans”: a list, a list of passages.
Please adhere to the following steps:
- 1. Carefully read the medical question and its answer.
- 2. Review the content of the provided document.
or verify the provided answer.
- 3. Identify and extract the passage from the document that directly supports the correct answer to the question.
- 4. If no passage in the document can directly support the correct answer or answer the question, return an empty list.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Here is the medical question, its answer, and the related document
Question: [QUESTION]
Answer: [ANSWER]
Document: [DOCUMENT]
Table 12. Prompts for Data Processing (2).
Answer by Evidence Prompt (LLM.answer)
You will be given a medical exam question and one or more evidence spans that were extracted from related documents. Your task is to
provide a detailed and comprehensive answer to the question based solely on the provided evidence spans.
Your output is a JSON object, which must contain the following keys:
- ”answer”: a string, the answer you derive from the reference documents.
- ”reason”: a detailed explanation of your reasoning process leading to the answer.
Please adhere to the following steps:
- 1. Review the exam question.
- 2. Review the provided evidence spans.
- 3. Based solely on the information contained in the evidence spans, provide a detailed and comprehensive answer to the question.
- 4. If the evidence spans do not provide sufficient information to answer the question, state ”The evidence passage can not answer the
question.” in ”answer” and explain why. If you don’t know the answer, don’t guess.
You must not use any common knowledge, personal knowledge, or external information beyond the provided evidence spans.
The ”answer” and ”reason” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the exam question and reference documents.
Question: [QUESTION]
Evidence Spans: [EVIDENCE SPANS]
Validate Answer Prompt (LLM.validate)
You will be given a medical question, a reference (standard) answer, and a model-generated answer. Your task is to evaluate the content
similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning.
Your output is a JSON object, which must contain the following keys:
- ”similarity_score”: a number between 0 and 1 indicating the content similarity between the two answers.
- ”explanation”: a detailed explanation of the similarities or differences that justify your similarity score.
Please adhere to the following steps:
- 1. Carefully read the medical question.
- 2. Review the reference answer and the model-generated answer.
- 3. Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion.
- 4. Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different.
- 5. Justify your score by explaining the similarities or differences between the two answers.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the question, standard answer, and generated answer.
Question: [QUESTION]
Reference Answer: [REFERENCE ANSWER]
Model-generated Answer: [MODEL-GENERATED ANSWER]
Query-Document Relevance Prompt (LLM.filter_score)
You will be given a medical search query and its associated passage. Your task is to evaluate the quality of query-passage pairs intended for
use in a medical encyclopedia knowledge retrieval evaluation dataset.
Your output is a JSON object, which must contain the following keys:
- ”quality_score”: an integer, a score from 1 to 5.
- ”explanation”: a string, providing a brief rationale for the given score.
Please adhere to the following steps:
- 1. Carefully read the query to understand the user’s information need.
- 2. Review the passage to assess its relevance and targeted content in relation to the query.
- 3. Assign a quality score from 1 to 5 and explain your reasoning.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the query and passage. Query: [QUERY]
Passage: [PASSAGE]
Figure 3. An overview of the diverse datasets in CMIRB benchmark.

### v2

@@ A.4 More Experiment Results
Table 11: Performance of different combinations of generators and retrievers on CMIRB benchmark.
Dataset legal_ legalbench_ legalbench_ Average
summar. contracts_qa lobbying
BGE 59.99 73.52 91.51 75.01
HyDE 58.95 74.82 92.78 75.52
SL-HyDE 63.50 75.10 93.15 77.25
Table 12: Performance of SL-HyDE in legal domain.
Appendix B CMIRB Datasets
Table 13: Dataset collection sources and quantity statistics.
B.1 Data Process
We curated a substantial dataset from various medical resources, as presented in Table 13, which details the source distribution and data volume. Our data preprocessing pipeline, depicted in Figuer 3 and Algorithm 1, employs prompt templates outlined in Figure 4 and Figure 5.
Figure 3: CMIRB benchmark construction pipeline.
Initially, we use ChatGPT [^8] https://openai.com/chatgpt to perform medical relevance detection on the texts, eliminating non-medical content (lines 3-8). Subsequently, ChatGPT assesses query-document relevance, filtering out low-relevance examples (lines 27-33). Our relevance assessment considers semantic alignment and the practical significance of data samples for their respective tasks, as highlighted in prompt LABEL:tab:data-prompts2.
For the MedExam and DuBaike datasets, the direct query-document signal isn’t initially provided. Both queries and documents in the MedExam dataset originate from Work Jin et al. (2021), where 100 randomly selected questions have corpus documents containing evidence sufficient to
Algorithm 1 Data Preprocessing Pipeline
3:// Step 1: Filter out medically irrelevant
4:for each query $q\in Q$, $d\in D$ do
5:     $med_{score}\leftarrow\mathrm{LLM}.\text{med\_score}(q/d)$
6:     if $med_{score}<threshold$ then
7:         Remove $q/d$
8:     end if
9:end for
10:// Step 2: Matching positive pairs
11:if query-document matching then
12:     for each query $q\in Q$ do
13:         // Retrieve top-k documents
14:         $D_{k}\leftarrow\text{BM25}(q,D)$
15:         $D_{k}\leftarrow\mathrm{LLM}.\text{reranking}(q;D_{k})$
16:         // Extract evidence snippets
17:         $E_{k}\leftarrow\mathrm{LLM}.\text{extract\_evidence}(q,D_{k})$
18:         // Generate answers
19:         $A_{k}\leftarrow\mathrm{LLM}.\text{answer}(q,E_{k})$
20:         for each document $d_{i}$ do
21:              if $\mathrm{LLM}.\text{validate}(a_{i},d_{i})$ then
22:                  Store $(q,d_{i})$
23:              end if
24:         end for
25:     end for
26:end if
27:// Step 3: Filter out pseudo-relevant pairs
28:for each matched pair $(q,d)$ do
29:     $rel_{score}\leftarrow\mathrm{LLM}.\text{filter\_score}(q,d)$
30:     if $rel_{score}<threshold$ then
31:         Remove $(q,d)$
32:     end if
33:end for
answer them, verified manually by the authors. In the DuBaike dataset, queries from Baidu Search and Baidu Zhidao often match the content distribution of Baidu Baike. These factors allow us to design a query-matching algorithm to locate the valuable document.
Through this iterative loop of self-ranking, evidence searching, answering, and verification, combined with ChatGPT’s advanced knowledge capabilities, we ensure high-quality, highly relevant query-document pairs.
B.2 Data Example
The datasets we constructed encompass various real-world medical scenarios, with examples from 10 different datasets illustrated in Table 14 and Table 15. Queries can take the form of a medical paper title, a patient’s symptom description, or an exam question. Corresponding documents include abstracts of medical papers, doctor-patient diagnostic conversations, and reference materials for exam questions.
Medical Relevance Prompt You will receive a question-answer pair from Baidu Search. Your task is to evaluate whether the Q&A is related to the medical field and output the result in JSON format. The JSON object must include the following keys: - "reason": a string explaining the reason for your judgment. - "label": an int, 0/1. Please adhere to the following steps: - If the content mentioned in the question and answer includes medical information and is related to the medical field, the label should be 1. - If most of the content in the question and answer is unrelated to the medical field, the label should be 0. You need to make a judgment and provide a reason. Please output the result as required, and do not output any other content. Here is the text: Question: [QUESTION] Answer: [ANSWER]
Passage Reranking Prompt You will be given a medical question, a reference (standard) answer, and a model-generated answer. Your task is to evaluate the content similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning. Your output is a JSON object, which must contain the following keys: - "similarity_score": a number between 0 and 1 indicating the content similarity between the two answers. - "explanation": a detailed explanation of the similarities or differences that justify your similarity score. Please adhere to the following steps: - 1. Carefully read the medical question. - 2. Review the reference answer and the model-generated answer. - 3. Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion. - 4. Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different. - 5. Justify your score by explaining the similarities or differences between the two answers. The "explanation" should be in Chinese. and your output must always be a JSON object, do not output anything else. Now here are the question, standard answer, and generated answer. Question: [QUESTION] Reference Answer: [REFERENCE ANSWER] Model-generated Answer: [MODEL-GENERATED ANSWER]
Evidence Extracting Prompt You will be given a medical question, its answer and a related document. Your task is to extract evidence spans from the document that directly or indirectly support the answer to the medical question. Your output is a JSON object, which must contain the following keys: - "evidence_spans": a list, a list of passages. Please adhere to the following steps: - 1. Carefully read the medical question and its answer. - 2. Review the content of the provided document. - 3. Identify and extract the passage from the document that directly supports the correct answer to the question. - 4. If no passage in the document can directly support the correct answer or answer the question, return an empty list. The "explanation" should be in Chinese. and your output must always be a JSON object, do not output anything else. Here is the medical question, its answer, and the related document Question: [QUESTION] Answer: [ANSWER] Document: [DOCUMENT]
Figure 4: Prompt for data processing (I).
Answer by Evidence Prompt You will be given a medical exam question and one or more evidence spans that were extracted from related documents. Your task is to provide a detailed and comprehensive answer to the question based solely on the provided evidence spans. Your output is a JSON object, which must contain the following keys: - "answer": a string, the answer you derive from the reference documents. - "reason": a detailed explanation of your reasoning process leading to the answer. Please adhere to the following steps: - 1. Review the exam question. - 2. Review the provided evidence spans. - 3. Based solely on the information contained in the evidence spans, provide a detailed and comprehensive answer to the question. - 4. If the evidence spans do not provide sufficient information to answer the question, state "The evidence passage can not answer the question." in "answer" and explain why. If you don’t know the answer, don’t guess. You must not use any common knowledge, personal knowledge, or external information beyond the provided evidence spans. The "answer" and "reason" should be in Chinese. and your output must always be a JSON object, do not output anything else. Now here are the exam question and reference documents. Question: [QUESTION] Evidence Spans: [EVIDENCE SPANS]
Validate Answer Prompt You will be given a medical question, a reference (standard) answer, and a model-generated answer. Your task is to evaluate the content similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning. Your output is a JSON object, which must contain the following keys: - "similarity_score": a number between 0 and 1 indicating the content similarity between the two answers. - "explanation": a detailed explanation of the similarities or differences that justify your similarity score. Please adhere to the following steps: - 1. Carefully read the medical question. - 2. Review the reference answer and the model-generated answer. - 3. Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion. - 4. Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different. - 5. Justify your score by explaining the similarities or differences between the two answers. The "explanation" should be in Chinese. and your output must always be a JSON object, do not output anything else. Now here are the question, standard answer, and generated answer. Question: [QUESTION] Reference Answer: [REFERENCE ANSWER] Model-generated Answer: [MODEL-GENERATED ANSWER]
Query-Document Relevance Prompt You will be given a medical search query and its associated passage. Your task is to evaluate the quality of query-passage pairs intended for use in a medical encyclopedia knowledge retrieval evaluation dataset. Your output is a JSON object, which must contain the following keys: - "quality_score": an integer, a score from 1 to 5. - "explanation": a string, providing a brief rationale for the given score. Please adhere to the following steps: - 1. Carefully read the query to understand the user’s information need. - 2. Review the passage to assess its relevance and targeted content in relation to the query. - 3. Assign a quality score from 1 to 5 and explain your reasoning. The "explanation" should be in Chinese. and your output must always be a JSON object, do not output anything else. Now here are the query and passage. Query: [QUERY] Passage: [PASSAGE]
Figure 5: Prompt for data processing (II).
MedExam
{CJK}UTF8gbsn Query: 问题: 胃癌最常发生的转移途径是()。选项: A:直接蔓延, B:血性转移, C:种植转移, D:淋巴转移, E:沿肠管转移.
(EN) Question: The most common metastasis route for gastric cancer is (). Options: A: Direct spread, B: Hematogenous metastasis, C: Seeding metastasis, D: Lymphatic metastasis, E: Along the intestinal tract.
{CJK}UTF8gbsn Document: 外科学 3.胃癌的扩散与转移 (2)淋巴转移:是胃癌的主要转移途径,进展期胃癌的淋巴转移率高达70%左右,侵及黏膜下层的早期胃癌淋巴转移率近20%。通常将引流胃的淋巴结分为16组,有的组还可以进一步分为若千亚组…
(EN) Surgery 3. Gastric cancer dissemination and metastasis (2) Lymphatic metastasis: It is the primary route of metastasis for gastric cancer, with a lymphatic metastasis rate of about 70% in advanced gastric cancer and approximately 20% in early gastric cancer invading the submucosa. Lymph nodes draining the stomach are usually classified into 16 groups, with some groups further divided into several subgroups…
DuBaike
{CJK}UTF8gbsn Query: 强迫症的表现是什么?
(EN) What are the manifestations of obsessive-compulsive disorder (OCD)?
{CJK}UTF8gbsn Document: 强迫症 临床表现 多发人群焦虑症与遗传因素、个性特点、不良事件、应激因素等均有关系,尤其与患者的个性特点紧密相关,比如:过分追求完美、犹豫不决、谨小慎微、固执等,具备这些不良个性特征容易患强迫症…
(EN) Obsessive-Compulsive Disorder Clinical Manifestations Prevalent Population Anxiety disorders are related to genetic factors, personality traits, adverse events, and stress factors, particularly closely linked to the patient’s personality traits. For instance, excessive perfectionism, indecisiveness, meticulousness, and stubbornness are traits that increase the risk of developing OCD…
DXYDisease
{CJK}UTF8gbsn Query: 维生素 A 缺乏症者需要做哪些检查来诊断?
(EN) What tests are needed to diagnose vitamin A deficiency?
{CJK}UTF8gbsn Document: 最准确的就是血液学检查。抽血检查血清维生素 A 的水平,对于成人来说,如果在 1.05~3.15 μmol/L,那么就表明不存在维生素 A 缺乏。 如果低于参考范围下限,那就是维生素 A 缺乏了。 …
(EN) The most accurate test is a hematological examination. A blood test to check the serum vitamin A levels is conducted. For adults, if the levels are between 1.05 and 3.15 umol/L, it indicates that there is no vitamin A deficiency. If the levels are below the lower limit of the reference range, it indicates vitamin A deficiency….
MedicalRetrieval
{CJK}UTF8gbsn Query: 一般宝宝的肚脐眼要多久愈合?
(EN) How long does it take for a baby’s belly button to heal?
{CJK}UTF8gbsn Document: 你好,宝宝的肚脐一般是1-2周左右会好的,时间长的也有一个月的,不过这个时候可能会有脐茸了。
(EN) Hello, a baby’s belly button generally heals in about 1 to 2 weeks, although it may take up to a month in some cases. During this time, there might also be umbilical granuloma.
CmedqaRetrieval
{CJK}UTF8gbsn Query: 甲状腺手术后多久可以干活?
(EN) How long after thyroid surgery can one return to work?
{CJK}UTF8gbsn Document: 皮肤的修复一般由两周左右就会不影响你的颈部活动了,至于皮下软组织以及肌肉组织的修复可能时间长一下,一般一个月后就不会有明显影响了,你就可以工作了。工作中注意不要劳累,调整好自己的情绪。
(EN) The skin usually heals in about two weeks, and you should no longer have restrictions on neck movement. However, the repair of subcutaneous soft tissue and muscle tissue may take longer. Generally, after about a month, there should be no significant impact, and you can return to work. During work, be sure to avoid overexertion and manage your emotions well.
Table 14: Data example in CMIRB (I).
DXYConsult
{CJK}UTF8gbsn Query: 症状及患病时长:感冒,鼻炎,失去嗅觉一周。就医及用药情况:未就医,自行服用泰诺。需要解答的问题:鼻炎,失去嗅觉怎么办
(EN) Symptoms and Duration of Illness: Cold, rhinitis, loss of smell for one week. Medical Consultation and Medication: No medical consultation, self-medicated with Tylenol. Questions Needing Answers: What to do about rhinitis and loss of smell?
{CJK}UTF8gbsn Document: 你好,如果近期有这种感冒的病史的话,就会导致出现嗅觉功能下降,建议在口服感冒药的技术上的话,用海盐水冲洗鼻腔,一天两次,鼻喷辅舒良或者内舒拿看看效果,如果分泌过多的话,可以口服桉柠蒎胶囊,每天三次每次一粒。
(EN) Hello, if there has been a recent history of cold symptoms, this can lead to decreased olfactory function. It is recommended to use saline nasal irrigation twice a day while taking cold medicine. You may also try nasal sprays like Budesonide or Fluticasone to see if they help. If there is excessive secretion, you can take Eucalyptus and Menthol capsules, three times a day, one capsule each time.
CovidRetrieval
{CJK}UTF8gbsn Query: 如何对待因履行工作职责感染新冠肺炎的医务人员?
(EN) How should healthcare workers who contract COVID-19 while fulfilling their duties be treated?
{CJK}UTF8gbsn Document: …为进一步加强疫情防控期间医务人员防护工作,切实保障医务人员身心健康,现将有关要求通知如下:一、高度重视医务人员防护工作做好医务人员防护工作,是预防和减少医务人员感染的关键举措,…
(EN) …To further enhance the protection of healthcare workers during the pandemic and ensure their physical and mental well-being, the following requirements are hereby notified: Pay great attention to the protection of healthcare workers Ensuring proper protection for healthcare workers is a key measure in preventing and reducing infections among them, …
IIYiPost
{CJK}UTF8gbsn Query: 病例讨论:静脉输入阿昔洛韦2天,出现腰痛、尿少
(EN) Case Discussion: Two days of intravenous acyclovir, followed by lower back pain and reduced urine output
{CJK}UTF8gbsn Document: 1.病例资料,患者,男,31岁。因静脉输入阿昔洛韦2天,出现腰痛、尿少伴恶心、呕吐6天入院。患者8天前因受凉感冒,出现咳嗽、发热(最高体温38.6℃),无明显咳痰,院外静脉给予NS500ml+青霉素钠盐800万U,vd,1次/日,…
(EN) Case Data, Patient: Male, 31 years old. The patient was admitted after experiencing lower back pain and reduced urine output, accompanied by nausea and vomiting for six days following two days of intravenous acyclovir administration. Eight days prior, the patient had caught a cold due to exposure, presenting with a cough and fever (highest temperature of 38.6°C), without significant sputum production. He received intravenous administration of …
CSLCite
{CJK}UTF8gbsn Query: 微球在组织工程中的应用
(EN) Application of Microspheres in Tissue Engineering
{CJK}UTF8gbsn Document: 背景:骨组织工程骨构建中如何使生长因子持续高效发挥作用是影响成骨速度和质量的关键,现多以各种材料的微球或支架作为缓释载体,但缓释作用有待提高.目的:实验拟制备壳聚糖微球,然后复合到纳米羟基磷灰石/聚乳酸羟基乙酸支架上…
(EN) Background: In bone tissue engineering, maintaining the sustained and efficient activity of growth factors is key to influencing the speed and quality of bone formation. Currently, microspheres or scaffolds made from various materials are commonly used as sustained-release carriers, but the release efficiency needs improvement. Objective: This experiment aims to prepare chitosan microspheres and incorporate them into a nano-hydroxyapatite/polylactic-co-glycolic acid (nHA/PLGA) scaffold, …
CSLRel
{CJK}UTF8gbsn Query: 高血压病的辨治及预防 高血压病可归属中医学"眩晕"、"头痛"等范畴,其起病隐匿,不易引起患者的充分重视,中后期可致心脑血管疾病、肾损害…
(EN) Differentiation and Treatment of Hypertension and Its Prevention Hypertension can be categorized under the terms "dizziness" and "headache" in traditional Chinese medicine (TCM). Its onset is insidious, often not receiving enough attention from patients, …
{CJK}UTF8gbsn Document: 辨证施治高血压 高血压病是现代医学病名,在中医归属眩晕病范畴,中医认为高血压与风、火、痰、虚有关.高血压的界定根据世界卫生组织(WHO)的标准,成人在休息状态下,收缩压持续高于140毫米汞柱…
(EN) TCM Syndrome Differentiation and Treatment of Hypertension Hypertension is a modern medical term, categorized under dizziness in TCM. TCM holds that hypertension is related to wind, fire, phlegm, and deficiency …
Table 15: Data example in CMIRB (II).